# Search for echoes on the edge of quantum black holes

### v2

@@ III Combination of Bayes factors assuming a target distribution
This section, instead of using the hierarchical combination of Bayes factors as in the former approach, focuses on extracting distribution parameters from multiple uncertain observations that are subject to selection biases. To achieve this, I assume a target Gaussian distribution $\bar{\mathcal{P}}(\rm{A},\bar{A},\sigma)=e^{-(A-\bar{A})^{2}/2\sigma^{2}}/\sqrt{2\pi\sigma^{2}}$ with a mean of $\bar{A}$ and a standard deviation of $\sigma$ for the measured parameter in this study. Then I combine them hierarchically Isi et al. (2019); Mandel et al. (2019),
@@
IV False positive and true positive estimation for GW190521
IV.1 Using samples from GR posterior
This is the same analysis performed in Abedi et al. (2021). In this section we perform the analysis using NRSur7dq4 waveform Varma et al. (2019). Thus, for consistency the Bayes factors of injections are compared to the one obtained with this waveform. Even though this section focuses on the outcome of the single event GW190521, it satisfies the validation test for ensuring the proper functioning of the method and model. Since in this section we don’t need to combine events and making a grid for amplitude A as the former section, I made it as free parameter inside its prior range of the model $0\leq A\leq 2$. In order to evaluate the true positive we have to inject GR models containing echoes. The GR only model for injections are drawn from random samples given by GR posterior, which is obtained from 10 different seed runs. Note that in the GR run, due to the marginalization over the coalescence phase of the binary, this particular parameter was not included in the corresponding posterior sample. As a result, the random samples consist of 14 parameters, and the coalescence phase was estimated additionally using waveform fitting to the data. Once the GR part is selected out of random samples of its posterior, the echo part is added. The echo part is within a random selection from uniform priors of $-13\leq\log_{10}\Lambda\leq 13$, $0\leq A\leq 2$, and $0\leq\phi\leq 2\pi$ which is given in first section. In order to evaluate the background and minimize non-Gaussian and/or non-stationary effects the data around the event is being used. Therefore, they are injected at random offset times of ±$\pm[5,32]$sec compared to GW190521 merger time. The gap ±$\pm 5$sec around GW190521 is due to preventing any contamination of data. In order to compare the foreground (true positive) to the background (false positive), I also injected a sample with zero echo amplitude for any corresponding GR+echo injection at the same offset time. In the next step run the same script job for injections with GR+echo and GR waveform. The number of injections performed for zero amplitude echoes (GR waveforms) is 343, while this number is 110 for non-zero amplitude echoes (GR+echo waveform). Fig. 7 is the results of this analysis in a histogram plot. We find injection Bayes factor for GR+echo with $34.5\pm 7.3\%$ and GR with $1.46^{+1.17}_{-0.88}\%$ being larger than $\mathcal{B}=7.5$ the echo Bayes factor for GW190521 with NRSur7dq4 waveform Abedi et al. (2021). Then the empirical likelihood ratio $=24.3^{+42.7}_{-11.8}$ which is the ratio of the foreground probability to the background probability is obtained. Please note that the errors associated with these estimates, as well as those in the subsequent subsection, are calculated using Bootstrapping.
Figure 7: Histograms to quantify false positive (GR injections) and true positive (GR+echo injections). Comparing to the GW190521 echo, we obtain their values as $1.46^{+1.17}_{-0.875}\%$, and $34.5\pm 7.3\%$ respectively.
IV.2 Using GW190521 signal
$\exp{(-\frac{(t-t_{\rm{merger}})^{2}}{\sigma^{2}})}\times\begin{pmatrix}\rm{Hanford\ data}\\ \rm{Livingston\ data}\\ \rm{Virgo\ data}\end{pmatrix}$ (6)
where $t_{\rm{merger}}=1242442967.4$ sec (GPS time). The parameter $\sigma$ should be taken in such that covers the main event signal and does’t leak to $\sim 1$ sec echo range. Thus $\sigma=0.5$ sec is chosen. Then instead of GR part in former section we replace it with this extracted signal.
After preparing the injections, we follow the same procedure as described earlier, running the script for injections using both the GR+echo and GR waveforms. A total of 250 injections are performed for zero-amplitude echoes (main event injections), while 20 injections are conducted for non-zero amplitude echoes (main event+echo injections). The results of this analysis are presented in Fig. 8 as a histogram plot. We observe that the injection Bayes factor for the main event+echo injections and for main event injections being larger than $\mathcal{B}=7.5$ (echo Bayes factor for GW190521 with the NRSur7dq4 waveform Abedi et al. (2021)) is $35\pm 15\%$, and $4.4^{+2.0}_{-2.0}\%$ respectively. Subsequently, we calculate the empirical likelihood ratio as $7.8^{+7.8}_{-4.0}$, representing the ratio of foreground probability to background probability.
V Conclusion and discussions
@@
The only event that has shown evidence for preference of GR+echo model is GW190521 with $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}=9.2$ (see Fig. 4). This is the most massive and confidently detected BBH merger event observed to date Abbott et al. (2021a). The remarkably loud ringdown of this event warrants special attention, making it a prime candidate for investigating gravitational wave echoes. After conducting a thorough analysis of this outlier event using two distinct methods, I found false detection probabilities of $1.46^{+1.17}_{-0.88}\%$ and $4.4^{+2.0}_{-2.0}\%$. The ratio of foreground probability to background in each case yielded an empirical likelihood ratio of $24.3^{+2.7}_{-11.8}$ and $7.8^{+7.8}_{-4.0}$. I refer the detailed interpretation and investigation about this event to Abedi et al. Abedi et al. (2021).
@@
It is worth to note that I didn’t see any evidence for echoes in O1 in contrast to Abedi et al. (2017a); Westerweck et al. (2018); Nielsen et al. (2019), possibly because the model I used here is different and has much suppressed amplitudes in contrast to ADA model in Abedi et al. (2017a).
It is crucial to underscore that, considering the present sensitivity levels of gravitational wave detectors Longo Micchi et al. (2021); Ma et al. (2022), there’s no concrete evidence of stimulated Hawking radiation from black hole mergers or Boltzmann echoes at $\geq 5\sigma$. Taking a cautious approach, this analysis stands as the most stringent constraint on the amplitude of this signal to date.

### v3

@@ 3.2 Hierarchical combination of Bayes factors assuming a target distribution
This section, instead of using the hierarchical combination of Bayes factors as in the former approach, focuses on extracting distribution parameters from multiple uncertain observations that are subject to selection biases. To achieve this, I assume a target Gaussian distribution $\bar{\mathcal{P}}(\rm{A},\bar{A},\sigma)=e^{-(A-\bar{A})^{2}/2\sigma^{2}}/\sqrt{2\pi\sigma^{2}}$ with a mean of $\bar{A}$ and a standard deviation of $\sigma$ for the measured parameter in this study. Then I combine them hierarchically [51, 52],
@@
4 Conclusion and discussion
@@
The only event that has shown evidence for preference of GR+echo model is GW190521 with $\mathcal{B}^{\rm{GR+echo}}_{\rm{GR}}=9.2$ (see Fig. 4). This is the most massive and confidently detected BBH merger event observed to date [5]. The remarkably loud ringdown of this event warrants special attention, making it a prime candidate for investigating gravitational wave echoes. After conducting a thorough analysis of this outlier event using two distinct methods, I found false detection probabilities of $1.46^{+1.17}_{-0.88}\%$ and $4.4^{+2.0}_{-2.0}\%$. The ratio of foreground probability to background in each case yielded an empirical likelihood ratio of $24.3^{+2.7}_{-11.8}$ and $7.8^{+7.8}_{-4.0}$. I refer the detailed interpretation and investigation about this event to Abedi et al. [26].
@@
It is worth to note that I didn’t see any evidence for echoes in O1 in contrast to [11, 16, 17], possibly because the model I used here is different and has much suppressed amplitudes in contrast to ADA model in [11].
It is crucial to underscore that, considering the present sensitivity levels of gravitational wave detectors [53, 54], there’s no concrete evidence of stimulated Hawking radiation from black hole mergers or Boltzmann echoes at $\geq 5\sigma$. Taking a cautious approach, this analysis stands as the most stringent constraint on the amplitude of this signal to date.